OpenStudy (anonymous):

one more , 14b-2=19b+33

OpenStudy (anonymous):

are u trying to find what b =

OpenStudy (anonymous):

yesss

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

b=-7

OpenStudy (anonymous):

Well I know that , I just need to find out how to solve that

OpenStudy (anonymous):

sorry i should have explained

OpenStudy (anonymous):

ow heres my work

OpenStudy (anonymous):

14b-2=19b+33 -14b -14b -2=5b+33 -33 -33 -35=5b /5 /5 -7=b

OpenStudy (anonymous):

thank you soo much , im going to study thiss